# Electronic structure and transport in materials with flat bands: 2D materials and quasicrystals

### v1

@@ 2.2.2 Sub-diffusive states in 2D quasiperiodic tilings
The specific role of long-range quasiperiodic order in electronic properties is still an open question in spite of a large number of studies (see, e.g., Refs. [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 35, 36, 58, 37, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68] and Refs. therein). Numerous studies confirm the existence of critical states, which are neither extended nor localized but are characterized by a power-law temporal decay of the wave function envelope at large distances. The existence of such critical states has been proven only for 1D quasiperiodic chains and 2D quasiperiodic tilings at certain very specific energies. In the presence of critical states, charge carrier diffusion at sufficiently large times $t$ follows a power law, the average spreading, $\Delta X=\sqrt{\Delta X^{2}}$, is expected to be [54, 55, 56],
@@
    For $10.5<\beta<$ 1, states are called super-diffusive.
@@
    For $\beta=0.5$ the regime is called diffusive. In disordered 2D systems, this regime always exists [189], but it may be reached for a very large time.
@@
    For $0.50<\beta<$ 0.5, states are called sub-diffusive.
@@
When disorder is introduced in the perfect approximant or perfect quasicrystal in the form of static defects (elastic scatterers) and/or inelastic scattering (temperature, magnetic field, …), the defects induce scattering and the propagation of the wave packet is expected to be diffusive for times $t$ larger than a scattering time $\tau$. The diffusivity $\mathcal{D}$ of charge carriers at energy $E$ can then be estimated by, [^1] This approach is less accurate than the Relaxation Time Approximation (RTA) we used to calculate conductivity in systems with defects (chapter 3 and A.1.2), but it gives qualitatively similar results.
$\mathcal{D}(E,\tau)\approx\frac{\Delta X^{2}(E,t=\tau)}{\tau}\propto\tau^{2\beta(E)-1},$ (13)
@@
The case $10.5<\beta<$ 1 (super-diffusive regime), leads to transport properties similar to a metal, since the conductivity decreases when disorder increases –i.e., when $\tau$ decreases–. Conversely, for $0.50<\beta<$ 0.5 (sub-diffusive regime), the conductivity increases when disorder increases as in real quasicrystals. Many authors consider [50, 51, 58, 54, 55, 56] that critical states could lead to $\beta<0.5$ but it has not yet been shown in 2D or 3D quasiperiodic structures (except for some very specific energies).
We have studied [60, 61, 62] two 2D quasiperidoc tilings, the octagonal –Ammann-Benenker– tiling [190, 191, 192, 193, 67] and the Penrose tiling [193, 194]. Our numerical calculations require the study of periodic structures (see A.1.3), so we have studied crystalline approximant structure series of increasingly large size to find the effect of long-range quasiperiodicity. The periodic approximants of the octagonal tiling were provided by C. Oguey according to [190], and those of the Penrose tiling by A. Szállás and A. Jagannathan according to [194]. Quantum diffusion in the smallest approximants, i.e., those with sufficiently small cell, has been studied from the eigenstates calculated by diagonalization (A.2). This allowed us to show [60, 61] the importance of non-Boltzmann terms in these systems as we found for the realistic approximants (Sec. 2.2.1). More recently, larger approximants have been studied by the recursion method in real space (A.3). Here, we focus exclusively on the latter approaches to discuss some general aspects of quantum diffusion in quasiperiodic tilings, specifically in the context of results obtained for octagonal tilings [62] and the Penrose tiling.
@@
Some results for octogonal approximants “with potential” [62] and Penrose approximants “pure hopping” are presented in Figs. 4 and 5, respectively. Note that qualitatively similar results are obtained for octogonal approximants “pure hopping” and Penrose approximants “with potential”.
@@
Figure 4: Density of states and quantum diffusion in octagonal approximants “with potential” (see text). (a) Series of periodic approximants proposed by M. Duneau et al. [190] indexed by $k$, with increasing unit cell size $L$. $N$ is the number of nodes per cell and $a$ the length of the edges. (b) Sketch of the tiling $k=4$. The blue dashed line is the cell. To simulate schematically a possible effect of the presence of different chemical elements, the on-site energies $\epsilon_{i}$ are proportional to the coordination of atoms (see text case “with potential”). (c) Total DOS $n$ in approximant $k=7$, due to recursion method $n$ is convoluted by a Lorentzian of width 5 meV. (d) Diffusion coefficient ${D}(t)=\Delta X(t)^{2}/t$ in approximant $k=9$ for various energies $E$. (e) ${D}(t)$ and (f) average spreading $\Delta X(t)$ at $E=2.7136t_{0}$ for different approximants. To have a realistic order of magnitude of time in seconds we used $t_{0}=1$ eV. The approximant tilings have been kindly provided by C. Oguey. Numerical method see A.3. Panels (c) and (d) are from [62].
@@
In the framework of the Kubo-Greenwood approach [30, 31, 32] for calculation of the conductivity, we use the polynomial expansion method developed by D. Mayou, S. N. Khanna, S. Roche, and F. Triozon [33, 34, 35, 36, 37] (see A.1.1 and A.3) to compute the mean square spreading $\Delta X(E,t)$ (Eq. (2)) of the wave packet at time $t$ and energy $E$, and the diffusion coefficient,
${D}(E,t)=\frac{\Delta X^{2}(E,t)}{t},$ (16)
@@
which tends towards diffusivity at large $t$ (A.1.1). ${D}(E,t)$ is shown in Figs. 4 and 5 for the octagonal and the Penrose tiling, respectively. The ballistic regime due to the periodicity of the approximant is reached at very large $t$, when $\Delta X(t)>L_{cell}$ where $L_{cell}$ is the approximant cell size; then $L(t)=V_{\mathrm{B}}t$, where $V_{\mathrm{B}}$ is the Boltzmann velocity, i.e., the intraband velocity in periodic approximants (Eq. (72) in the appendix). For the approximants with the largest cell and at realistic time range, Figs. 4 and  5 show that this Boltzmann term is negligible and, for all purposes of this discussion, the largest approximants are similar to the quasiperiodic system. For approximants with a smaller unit cell, at very large times, a ballistic behavior is clearly achieved which is due to the periodicity of the approximants. Indeed, we have checked that at these times, the mean spreading $\Delta X$ is larger than the cell size of the approximants. For the largest approximants, this regime has not yet been reached.
@@
Figure 5: Density of states and quantum diffustion in approximants of 2D Penrose tiling with “pure hopping” without defects (except dashed line of panel (e)): (a) Series of periodic approximants proposed by M. Duneau and M. Audier [194] indexed by an integer $T$, with increasing unit cell size $L$. $N$ is the number of nodes per cell and $a$ the size of the edges. (b) Sketch of a cell of the approximant $T=3$. For pure hopping Hamiltonian ($\epsilon_{i}=0$ and $t_{0}=1$ eV): (c) Total DOS of the approximant $T=3$ [195, 61]. (d) Diffusion coefficient ${D}(t)=\Delta X^{2}/t$ for a series of approximants at the same energy $E=0.047$ eV. See text for the definition of $L_{1}$ and $L_{2}$. (e) Effect on ${D}(t)$ of static defects in large approximant ($T=9$): ${D}(t)$ for two energies corresponding to sub-diffusive ($\beta<0.5$) and super-diffusive ($\beta>0.5$) states, respectively; (solid line) without defects, and (dashed line) with random distribution of Anderson defects $\epsilon_{i}\in[-W/2;W/2]$. To have a realistic order of magnitude of time in second we used the hopping term $t_{0}=1$ eV. The atomic positions of the approximants have been kindly provided by A. Szállás and A. Jagannathan [196, 197]. For the numerical method see A.3. Panel (c) is from [61].
@@
    At very small times, typically when $\Delta X(t)<a$ where $a$ is the length of the base tile edge, the mean spreading increases linearly with $t$, $\Delta X(t)=V_{0}t$ (ballistic behavior), where $V_{0}>V_{\mathrm{B}}$ [27, 158]. Indeed, the effective velocity $V_{0}$ includes intraband terms of the velocity operator and interband terms of the velocity operator between band states with small energy difference $\Delta E\lesssim\hbar/t$.
@@
    Another distance $L_{2}$ (respectively another time $t_{2}$, $L(t_{2})=L_{2}$) appears. For $\Delta X(t)>L_{2}\approx{\rm a~{}few~{}}10a$, a new regime appears and ${D}(t)$ follows a power law. It is thus characteristic of the medium and long-range quasiperiodic order. Figures 4(d) and 5(e) shows that the $\beta$ value can switch from a sub-diffusive regime ($\beta<0.5$) to a super-diffusive regime ($\beta>0.5$) over a small variation of energy. The $t_{2}$ values, $t_{2}\approx 10^{-13}$${10^{-14}}$ s, have the order of magnitude of the scattering time above which measurements show unusual transport properties in quasicrystals [41].
[…]
@@
Since the sub-diffusive states and the super-diffusive states are due to long-range quasiperiodic order their diffusive properties should be very sensitive to disorder. This is verified by the ${D}(E,t)$ calculation in the presence of a weak disorder (Anderson disorder) showing (Fig. 5(e)) that a sub-diffusive state and a super-diffusive state become diffusive in the presence of disorder. These numerical calculations clearly show that on one hand, at certain energies, when the state is sub-diffusive, the quasiperiodic order localizes the charge carriers sufficiently for their behavior to be similar to that of a semiconductor, and then the increase of disorder that destroys the effect of quasiperiodic order leads to an increase in diffusivity and thus in conductivity. This regime is therefore a semiconductor like regime. On the other hand, for other energies, when the state is super-diffusive, the regime is metallic like and the increase of disorder leads to a decrease in conductivity.
@@
To summarize, in the octagonal tiling and the Penrose tiling, the charge carrier propagation is determined by the wave-packet spreading in the quasiperiodic lattice. From numerical calculations, two length scales seem to characterize this quasiperiodic spreading. $L_{1}$, typically $L_{1}={\rm a~{}few~{}}a$, above which the propagation is almost diffusive in spite of the absence of static defects. $L_{2}$, typically $L_{2}={\rm a~{}few~{}}10a$, above which specific quasiperiodic symmetries lead to a power-law dependence of the root mean square spreading, $\Delta X(t)\propto t^{\beta}$. For some energies, states are super-diffusive or diffusive, i.e., $\beta\geq 0.5$, whereas for other energies, a sub-diffusive regime, i.e., $\beta<0.5$, sets in as expected for critical states characteristic of quasiperiodicity. This sub-diffusive regime is the generalization to quasicrystals of the non-Boltzmann propagation (section 2.1) found in realistic approximants of i-AlMnSi and i-AlCuFe, the complex intermetallic alloys $\lambda$-AlMn (section 2.2.1), and in small approximants of octagonal and Penrose tilings [60, 61]. This sub-diffusive regime, which is not created by disorder in the atomic structure, but by the quasiperiodic order itself, results in a particular localization that corresponds to the SVR and that can be destroyed by disorder, resulting in semiconductor like behavior.
@@ 2.3 Twisted bilayer graphene (TBG)
Experimentally, graphene can be formed in multilayers on SiC [198, 199, 200, 201, 202, 203, 204] but also on metal surfaces such as Ni [205] and in exfoliated flakes [96] where hopping terms between successive layers play a crucial role. While on the Si face of SiC, multilayers have an AB Bernal stacking with electronic properties different from graphene [206, 198, 199, 207, 208, 135], on the C-face multilayers are twisted multilayers of graphene with various angles of rotation between two successive layers. For large rotation angle $\theta$ between two layers, multilayers show graphene-like properties even when they involve a large number of graphene layers. Indeed, as shown by ARPES [202, 203, 204], STM [209, 97], transport [75], and optical transitions [76], their properties are characteristic of a linear graphene-like dispersion at K points of the Brillouin zone (Dirac cone). Therefore, in TBG interlayer hopping terms do not systematically destroy graphene-like properties, but they can lead to the emergence of new behavior induced by the moiré patterns [98, 99, 100] (Fig. 6(a)) that is accentuated when the angle $\theta$ is very small [89]. Since the early 2010s, theoretical studies [90, 91, 92, 93] and experimental evidences [96, 97] have shown that moiré TBGs can confine conduction electrons in a controlled manner.

### v2

@@ 2.2.2 Sub-diffusive states in 2D quasiperiodic tilings
The specific role of long-range quasiperiodic order in electronic properties is still an open question in spite of a large number of studies (see, e.g., Refs. Kohmoto86 , Fujiwara89_Fibo , Sire90 , Passaro92 , Sire94 , Mayou94 , Guarneri94 , Piechon96 , SchulzBaldes98 , Zhong98 , Roche97 , Roche99 , Mayou00 , Triozon02 , Jagannathan07 , Trambly11_QC , Trambly14_QC , Trambly17 , Collins17 , Chiaracane21 , Grimm21 , Jagannathan23 , Jagannathan24 , Bellissard24 and Refs. therein). Numerous studies confirm the existence of critical states, which are neither extended nor localized but are characterized by a power-law temporal decay of the wave function envelope at large distances. The existence of such critical states has been proven only for 1D quasiperiodic chains and 2D quasiperiodic tilings at certain very specific energies. In the presence of critical states, charge carrier diffusion at sufficiently large times $t$ follows a power law, the average spreading, $\Delta X=\sqrt{\Delta X^{2}}$, is expected to be Guarneri94 , Piechon96 , SchulzBaldes98 ,
@@
    For $0.5<\beta<1$, states are called super-diffusive.
@@
    For $\beta=0.5$ the regime is called diffusive. In disordered 2D systems, this regime always exists Lee85 , but it may be reached for a very large time.
@@
    For $0<\beta<0.5$, states are called sub-diffusive.
@@
When disorder is introduced in the perfect approximant or perfect quasicrystal in the form of static defects (elastic scatterers) and/or inelastic scattering (temperature, magnetic field, …), the defects induce scattering and the propagation of the wave packet is expected to be diffusive for times $t$ larger than a scattering time $\tau$. The diffusivity $\mathcal{D}$ of charge carriers at energy $E$ can then be estimated by, [^1] This approach is less accurate than the Relaxation Time Approximation (RTA) we used to calculate conductivity in systems with defects (section 3 and A.1.2), but it gives qualitatively similar results.
$\mathcal{D}(E,\tau)\approx\frac{\Delta X^{2}(E,t=\tau)}{2\tau}\propto\tau^{2\beta(E)-1},$ (13)
@@
The case $0.5<\beta<1$ (super-diffusive regime), leads to transport properties similar to a metal, since the conductivity decreases when disorder increases –i.e., when $\tau$ decreases–. Conversely, for $0<\beta<0.5$ (sub-diffusive regime), the conductivity increases when disorder increases as in real quasicrystals. Many authors consider Sire90 , Passaro92 , Mayou00 , Guarneri94 , Piechon96 , SchulzBaldes98 that critical states could lead to $\beta<0.5$ but it has not yet been shown in 2D or 3D quasiperiodic structures (except for some very specific energies).
We have studied Trambly11_QC , Trambly14_QC , Trambly17 two 2D quasiperidoc tilings, the octagonal –Ammann-Benenker– tiling Duneau89 , Socolar89 , Ammann92 , Katz94 , Jagannathan24 and the Penrose tiling Katz94 , Duneau94 . Our numerical calculations require the study of periodic structures (see A.1.4), so we have studied crystalline approximant structure series of increasingly large size to find the effect of long-range quasiperiodicity. The periodic approximants of the octagonal tiling were provided by C. Oguey according to Duneau89 , and those of the Penrose tiling by A. Szállás and A. Jagannathan according to Duneau94 . Quantum diffusion in the smallest approximants, i.e., those with sufficiently small cell, has been studied from the eigenstates calculated by diagonalization (A.2). This allowed us to show Trambly11_QC , Trambly14_QC the importance of non-Boltzmann terms in these systems as we found for the realistic approximants (Sec. 2.2.1). More recently, larger approximants have been studied by the recursion method in real space (A.3). Here, we focus exclusively on the latter approaches to discuss some general aspects of quantum diffusion in quasiperiodic tilings, specifically in the context of results obtained for octagonal tilings Trambly17 and the Penrose tiling.
@@
Some results for octogonal approximants “with potential” Trambly17 and Penrose approximants “pure hopping” are presented in Figs. 4 and 5, respectively. Note that qualitatively similar results are obtained for octogonal approximants “pure hopping” and Penrose approximants “with potential”.
@@
Figure 4: Density of states and quantum diffusion in octagonal approximants “with potential” (see text). (a) Series of periodic approximants proposed by M. Duneau et al. Duneau89 indexed by $k$, with increasing unit cell size $L$. NN is the number of nodes per cell and $a$ the length of the edges. (b) Sketch of the tiling $k=4$. The blue dashed line is the cell. To simulate schematically a possible effect of the presence of different chemical elements, the on-site energies $\epsilon_{i}$ are proportional to the coordination of atoms (see text case “with potential”). (c) Total DOS $n$ in approximant $k=7$, due to recursion method $n$ is convoluted by a Lorentzian of width 5 meV. (d) Diffusion coefficient $2{D}(t)=\Delta X(t)^{2}/t$ in approximant $k=9$ for various energies $E$. (e) 2D(t)2{D}(t) and (f) average spreading $\Delta X(t)$ at $E=2.7136t_{0}$ for different approximants. To have a realistic order of magnitude of time in seconds we used $t_{0}=1$ eV. The approximant tilings have been kindly provided by C. Oguey. For the numerical method see A.3. Panels (c) and (d) are from Trambly17 .
@@
In the framework of the Kubo-Greenwood approach Kubo57 , Greenwood58 , Kubo91 for calculation of the conductivity, we use the polynomial expansion method developed by D. Mayou, S. N. Khanna, S. Roche, and F. Triozon Mayou88 , Mayou95 , Roche97 , Roche99 , Triozon02 (see A.1.1 and A.3) to compute the mean square spreading $\Delta X(E,t)$ (Eq. (2)) of the wave packet at time $t$ and energy $E$, and the diffusion coefficient,
${D}(E,t)=\frac{\Delta X^{2}(E,t)}{2t},$ (16)
@@
which tends towards diffusivity at large $t$ (A.1.1). D(E,t){D}(E,t) is shown in Figs. 4 and 5 for the octagonal and the Penrose tiling, respectively. The ballistic regime due to the periodicity of the approximant is reached at very large $t$, when $\Delta X(t)>L_{cell}$ where $L_{cell}$ is the approximant cell size; then $L(t)=V_{\mathrm{B}}t$, where $V_{\mathrm{B}}$ is the Boltzmann velocity, i.e., the intraband velocity in periodic approximants (Eq. (76) in the appendix). For the approximants with the largest cell and at realistic time range, Figs. 4 and  5 show that this Boltzmann term is negligible and, for all purposes of this discussion, the largest approximants are similar to the quasiperiodic system. For approximants with a smaller unit cell, at very large times, a ballistic behavior is clearly achieved which is due to the periodicity of the approximants. Indeed, we have checked that at these times, the mean spreading $\Delta X$ is larger than the cell size of the approximants. For the largest approximants, this regime has not yet been reached.
@@
Figure 5: Density of states and quantum diffustion in approximants of 2D Penrose tiling with “pure hopping” without defects (except dashed line of panel (e)): (a) Series of periodic approximants proposed by M. Duneau and M. Audier Duneau94 indexed by an integer $T$, with increasing unit cell size $L$. NN is the number of nodes per cell and $a$ the size of the edges. (b) Sketch of a cell of the approximant $T=3$. For pure hopping Hamiltonian ($\epsilon_{i}=0$ and $t_{0}=1$ eV): (c) Total DOS of the approximant $T=3$ Zijlstra00 , Trambly14_QC . (d) Diffusion coefficient $2{D}(t)=\Delta X^{2}/t$ for a series of approximants at the same energy $E=0.047$ eV. See text for the definition of $L_{1}$ and $L_{2}$. (e) Effect on 2D(t)2{D}(t) of static defects in large approximant ($T=9$): D(t){D}(t) for two energies corresponding to sub-diffusive ($\beta<0.5$) and super-diffusive ($\beta>0.5$) states, respectively; (solid line) without defects, and (dashed line) with random distribution of Anderson defects $\epsilon_{i}\in[-W/2;W/2]$. To have a realistic order of magnitude of time in second we used the hopping term $t_{0}=1$ eV. The atomic positions of the approximants have been kindly provided by A. Szállás and A. Jagannathan Jagannathan07_penroseAF , Szallas08 . For the numerical method see A.3. Panel (c) is from Trambly14_QC .
@@
    At very small times, typically when $\Delta X(t)<a$ where $a$ is the length of the base tile edge, the mean spreading increases linearly with $t$, $\Delta X(t)=V_{0}t$ (ballistic behavior), where $V_{0}>V_{\mathrm{B}}$ Trambly06 , Trambly14_CRAS . Indeed, the effective velocity $V_{0}$ includes intraband terms of the velocity operator and interband terms of the velocity operator between band states with small energy difference $\Delta E\lesssim\hbar/t$.
@@
    Another distance $L_{2}$ (respectively another time $t_{2}$, $\Delta X(t_{2})=L_{2}$) appears. For $\Delta X(t)>L_{2}\approx{\rm a~few~}10a$, a new regime appears and D(t){D}(t) follows a power law. It is thus characteristic of the medium and long-range quasiperiodic order. Figures 4(d) and 5(e) shows that the $\beta$ value can switch from a sub-diffusive regime ($\beta<0.5$) to a super-diffusive regime ($\beta>0.5$) over a small variation of energy. The $t_{2}$ values, $t_{2}\approx 10^{-13}$$10^{-14}$} s, have the order of magnitude of the scattering time above which measurements show unusual transport properties in quasicrystals Berger94 .
Both distances $L_{1}$ (a few $a$), $L_{2}$ (a few 10aa), and the exponent $\beta$ at time $t>t_{2}$, depend a lot on the energy value $E$. $L_{1}<L_{2}$, but at some energy it even seems that $L_{1}\approx L_{2}$. Further analysis is necessary to understand this energy dependence.
@@
Since the sub-diffusive states and the super-diffusive states are due to long-range quasiperiodic order their diffusive properties should be very sensitive to disorder. This is verified by the D(E,t){D}(E,t) calculation in the presence of a weak disorder (Anderson disorder) showing (Fig. 5(e)) that a sub-diffusive state and a super-diffusive state become diffusive in the presence of disorder. These numerical calculations clearly show that on one hand, at certain energies, when the state is sub-diffusive, the quasiperiodic order localizes the charge carriers sufficiently for their behavior to be similar to that of a semiconductor, and then the increase of disorder that destroys the effect of quasiperiodic order leads to an increase in diffusivity and thus in conductivity. This regime is therefore a semiconductor like regime. On the other hand, for other energies, when the state is super-diffusive, the regime is metallic like and the increase of disorder leads to a decrease in conductivity.
@@
To summarize, in the octagonal tiling and the Penrose tiling, the charge carrier propagation is determined by the wave-packet spreading in the quasiperiodic lattice. From numerical calculations, two length scales seem to characterize this quasiperiodic spreading. $L_{1}$, typically $L_{1}={\rm a~few~}a$, above which the propagation is almost diffusive in spite of the absence of static defects. $L_{2}$, typically $L_{2}={\rm a~few~}10a$, above which specific quasiperiodic symmetries lead to a power-law dependence of the root mean square spreading, $\Delta X(t)\propto t^{\beta}$. For some energies, states are super-diffusive or diffusive, i.e., $\beta\geq 0.5$, whereas for other energies, a sub-diffusive regime, i.e., $\beta<0.5$, sets in as expected for critical states characteristic of quasiperiodicity. This sub-diffusive regime is the generalization to quasicrystals of the non-Boltzmann propagation (section 2.1) found in realistic approximants of i-AlMnSi and i-AlCuFe, the complex intermetallic alloys $\lambda$-AlMn (section 2.2.1), and in small approximants of octagonal and Penrose tilings Trambly11_QC , Trambly14_QC . This sub-diffusive regime, which is not created by disorder in the atomic structure, but by the quasiperiodic order itself, results in a particular localization that corresponds to the SVR and that can be destroyed by disorder, resulting in semiconductor like behavior.
@@ 2.3 Twisted bilayer graphene (TBG)
Experimentally, graphene can be formed in multilayers on SiC Ohta06 , Brihuega08 , Hass06 , Hass08b , Hass08 , Sprinkle09 , Hicks11 but also on metal surfaces such as Ni Luican11 and in exfoliated flakes Li10 where hopping terms between successive layers play a crucial role. While on the Si face of SiC, multilayers have an AB Bernal stacking with electronic properties different from graphene Latil06 , Ohta06 , Brihuega08 , Varchon08 , McCann13 , Rozhkov16 , on the C-face multilayers are twisted multilayers of graphene with various angles of rotation between two successive layers. For large rotation angle $\theta$ between two layers, multilayers show graphene-like properties even when they involve a large number of graphene layers. Indeed, as shown by ARPES Hass08 , Sprinkle09 , Hicks11 , STM Miller09 , Brihuega12 , transport Berger06 , and optical transitions Sadowski06 , their properties are characteristic of a linear graphene-like dispersion at K points of the Brillouin zone (Dirac cone). Therefore, in TBG interlayer hopping terms do not systematically destroy graphene-like properties, but they can lead to the emergence of new behavior induced by the moiré patterns Campanera07 , Mele10 , Gratias23 (Fig. 6(a)) that is accentuated when the angle $\theta$ is very small LopesdosSantos07 . Since the early 2010s, theoretical studies Trambly10 , Bistritzer10 , SuarezMorell10 , Bistritzer11 and experimental evidences Li10 , Brihuega12 have shown that moiré TBGs can confine conduction electrons in a controlled manner.